OpenStudy (anonymous):

what is the slop and y-intercept of 3x+4y=12

OpenStudy (anonymous):

3x+4y=12 -3x -3x 4y=-3x+12 ____________ 4 y=-(3/4)x+3 so slope is -(3/4) and y-intercept is 3